Ticket Number: HEA-NOC/20080331-2 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20080331 17:41 UTC+1 Problem Start: 20080401 11:00 UTC+1
Ticket Update: 20080402 15:49 UTC+1 Problem End: 20080401 17:00 UTC+1
Migration of NCRI's current connections to HEAnet to new IP Network
Service interruption is expected due to NCRI moving to another building,
this is covered in a separate ticket (#20080327-5). BGP sessions need to
be reset and
will be done gradually. A side effect of this migration is asymmetrical
routing while convergence takes place.
Once the changes are completed, traffic will use the new connections,
continuing to have the current ones as backup. Eventually these will be
Contingency plan is to revert changes to existing setup by disabling any
routing over the new links.
20080331 17:37 GL TG has confirmed with NCRI of date and time of move. TG
is currently in NCRI ready to make the necessary changes for this
migration to take place. Opening ticket
20080401 17:21 GL TG installed devices and migration took place and was
completed. NCRI's traffic is now routing via Citywest CRS-1 with
Kilcarbery CRS-1 acting a secondary route for redundancy. Will monitor for
20080402 15:49 GL Links stable. Closing ticket
This ticket can be monitored at http://www.hea.net/tickets/20080331-2
HEAnet Limited. Registered in Ireland, No. 275301.